In 3 years Darla will be twice as old as she was 8 years ago. How old is Darla today?
Answer by Maths68(1474)

Let
Present age of Darla = d


Her age in three years = d+3
Her age 8 years ago = d-8
In 3 years Darla will be twice as old as she was 8 years ago.
d+3=2(d-8)
d+3=2d-16
d-2d=-16-3
-d=-19
Multiply by -1 both sides
d=19
Darla is 19 years old Now!!!